The elemental analysis of an organic compound with general formula CxHYOZ of molar mass 130 gives the following results; C=64.5%, H=10.7%. The exact formula of this compound is?



1
Expert's answer
2021-09-02T10:18:11-0400

CXHYOz contains elements C, H and O with

% abundance 64.5%, 10.7% and 24.8% respectively.

Moles by percentage abundance; C= 64.5/12, H=10.7/1, O=24.8/16

C=5.375, H=10.7, O=1.55

Mole ratio after division using the smallest value; C:H:O= 3.5:7:1

Simple ratio derived from multiplying by 2 = 7:14:2

Emperical formula= C7H14O2

Molecular mass = (EMF) n

(C7H14O2 ) n =130

{(12x7) + (14x1) +(16x2)) n=130

130n=130; n=1

Hence formula C7H14O2